OpenStudy (anonymous):

How to find the x-intercept of x^2-5x+4=y

myininaya (myininaya):

let y=0 solve for x

myininaya (myininaya):

x^2-5x+4=0 (x-4)(x-1)=0 x=4 x=1

OpenStudy (anonymous):

wow..

OpenStudy (anonymous):

thanx for your help

OpenStudy (anonymous):

For problems you can't factor use the quadratic formula. Given \[ax^2+bx^2+c=0\]solve using\[x=[-b \pm \sqrt{b^2-4ac}]/2a\] The quadratic formula will always give you the answer (if one exists. i.e., if there are any x-intercepts). So for your problem we have\[x=[-(-5)\pm \sqrt{(-5)^2-4(1)(4)}]/2(1)\]or \[x=[5\pm \sqrt{25-16}]/2\] so \[x=4 \] and \[x=1\]